Micah is 73 years old. Alice is 31 years old. How many years ago was Micah's age 3 times Alice's age?

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52794)

.

73 - x = 3*(31-x)


73 - x = 93 - 3x


3x - x = 93 - 73


2x     = 20


x      = 20/2 = 10.


ANSWER.  10 years ago.

Solved.


Answer by greenestamps(13200)



Tutor @ikleyn has shown a perfectly good formal algebraic solution to the problem.

If formal algebra is not required, here is a quick and easy informal path to the solution.

(1) The difference between their ages is now (and always will be) 73-31 = 42 years.

(2) When Micah was 3 times Alice's age, the difference in their ages was twice Alice's age. (The difference between x and 3x is 2x). Solving, 2x = 42; x = 21; it was when Alice was 21 and Micah was 21+42 = 63 that Micah was 3 times as old as Alice.

(3) Since their ages are now 31 and 73, that was 10 years ago.

ANSWER: 10 years ago